Greenhab Report – February 10th

Crew 205 Green Hab Report – 10-FEB-2019

GreenHab Officer: Nathan Hadland

Environmental control: Heating and Cooling w/ ambient air (4 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 17.6°C

High: 27.8°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 11 gallons

Water in Blue Tank: ~ 148 gallons (NOTE – GreenHab tank will be filled later tonight by Atila)

Time(s) of watering for crops:

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)
[AT TIME OF WATERING]

1) 08.38 AM 18.0° C / 35 %
2) 13.37 PM 25.7° C / 11 %
3) 16.45 PM 21.1° C / 31 %
4) 19:00 PM 18.8° C / 42 %

Changes to crops: None

Narrative: The carrots appear to be dominating their portion of the planter and seem to be drawing nutrients from the adjacent oregano and parsley. Tuesday, we will harvest some of these carrots to help alleviate the stress on the rest of the grow bed and provide supplemental nutrition to the crew. The cucumber plants are dead and we will recover them and repurpose the soil for the incoming tomato plants at the beginning of sim tomorrow as well. We changed the fly paper in all of the pots and grow beds, and we took photos of all of the plants in the GreenHab to compare when we photograph them again several times throughout the sim.

Harvest: None

Support/supplies needed: I suggest purchasing packages of fly paper in the next few weeks, there is only one left.

Crew 204 Greenhab Report 09Feb2019

Crew 204 Green Hab Report – 9-FEB-2019

GreenHab Officer: Patil (Cosmos)

Environmental control: Heating and Cooling w/ ambient air (5 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 18.7°C

High: 25.6°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 12 gallons

Water in Blue Tank: ~ 159 gallons

Time(s) of watering for crops:

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)
[AT TIME OF WATERING]

1) 10.00 AM 18.1° C / 75 %
2) 13.22 PM 19.3° C / 36 %
3) 15:12 PM 23.4° C / 23 %
4) 20:40 PM 19.2° C / 49 %

Changes to crops: None

Narrative: We are having change over today, I am feeling so emotional to leave my kids(plants) in the hands of a new green officer Mr. Nathan Harland. I am sure he will take care of them and make them grow even better. Plants are growing as expected and looking forward to seeing their status in near future. Thank you MDRS for this amazing experience. This is going to be a life long experience. Thank you for everything.

Harvest: None

Support/supplies needed: None

Commander Report – February 9th

Author: Avishek Ghosh

Title: The words of wisdom.

Crew 204 Commander Report 09-FEB-2019.

The day had been spent cleaning the hab and finishing the incomplete work by the crew. We had lunch together made by Sonal (Mars_Bag) and all of us. This was followed by training given to Crew 205.

It is really hard to believe that the simulation has ended. We feel fortunate enough to get this opportunity to be a part of this simulation. This experience has been a worthy one as we learnt to face a lot of challenges and difficulties with a success in the mission.

The journey taught us cultural exchange though we belong to the same country. I believe this isolation will act as an influential factor in our lives. From the bottom of my heart, I am thankful to mission support, Mr. Atila and Mr. David for supporting us the fullest.

This is Avishek Ghosh, signing off Crew 204 from Mars Desert Research Station and welcoming Crew 205 and wishing them very good luck and best wishes for the mission.

Operations Report – February 10th

Crew 205 Operations Report 10 – FEB – 2019

SOL: 0

Name of person filing report: Veronica Trivino

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): Turned on this evening by 1800 (10thFeb), Stopped by this morning 0800 (10thFeb)

Solar— SOC % – Turned off Gen (10th Feb) 70%, Turned on Gen (10th Feb) 20%

Diesel Reading – 25%

Propane Reading – ~ 50%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 90%, ~ 495 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 148 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01410586 units

Toilet tank emptied: yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Used

Hours: 49 hours

Beginning charge: 100%

Ending charge: 88%

Currently charging: Yes

Curiosity rover used: Used

Hours: 73.6 hours

Beginning charge: 100%

Ending charge: 63%

Currently charging: Yes

Spirit rover used: Used

Hours: 73

Beginning charge: 100%

Ending charge: 85%

Currently charging: Yes

Notes on rovers: None

ATV’s Used: None

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Driven to Hanksville 4 times to get water

Crew Car used and why, where? Not used

General notes and comments: Hab Car is operational and back from workshop

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: None

Operations Report – February 9th

Crew 204 Operations Report 09 – FEB – 2019

SOL: 12

Name of person filing report: Kunal Naik (Space_bot)

Non-nominal systems: Generator

Notes on non-nominal systems: Generator: Non- nominal. The alternator needs to be replaced and the coolant system flushed and hoses replaced. All parts already arrived.

Generator (hours run): Turned on this evening by 1700 (9th Feb), Stopped by this morning 0830 (9th Feb)

Solar— SOC % – Turned off Gen (9th Feb) 72%, Turned on Gen (9th Feb) 55%

Diesel Reading – 28%

Propane Reading – ~ 50%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 50%, ~ 250 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 159 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01409959 units

Toilet tank emptied: yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Not Used

Hours: 48.5 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not Used

Hours: 73.6 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Not used

Hours: 72.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Curiosity returned yesterday with a very low charge, compared with Oppy. For long distance EVA., it’s recommended to use Oppy.

ATV’s Used: None

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: Nope

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? To town for Dinner

Crew Car used and why, where? To town for Dinner

General notes and comments: HabCar: back on campus. The car was taken to the workshop in order to fix the transmission issue. It was successfully repaired and ready to work. Currently, it doesn’t have a spare tire. The tire is on its way. One of the rear tires had a nail on it, it’s already patched.

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: None

Sol Summary – February 10th

Crew 205 Sol Summary Report 10-Jan-2019

Sol: 0

Summary Title: Ready for Mars!

Author’s name: Natalia Larrea

Mission Status: Nominal

Sol Activity Summary: Today was a special day: the first day we woke up at the Hab! We spent the night together with Crew 204. The place was crowded but we had a great evening conversing and playing board games together.

We woke up full of energy and ready to prepare the final details for our mission. In the morning, we got our first pictures as a crew and proceeded execute crew handover with Crew 204. We came back to the Hab and settled in to our new home for the next two weeks. We enjoyed blueberry pancakes for breakfast, unpacked our bags, did some deep cleaning, took inventory, and loaded the water tanks.

In the afternoon, Atila and David joined us at the Hab to conduct training on mission protocols and operations. Then we split into two groups for EVA training using the rovers in the field. Finally, we practiced putting on the suits and using the earpieces.

Look Ahead Plan: Tomorrow we enter sim. We will start our new daily routine and plan to conduct our first EVAs. Stay tuned for our future reports. We have two weeks of full of exciting projects coming up! We are ready for Mars!

Anomalies in work: None

Weather: Clear skies all day long, intermittent high intensity winds.

Crew Physical Status: Nominal

EVA: None

Reports to be filed: Sol Summary, EVA request, Operations Report, Greenhab Report, Journalist Report, HSO checklist, Crew Bios and pictures, mission patch

Support Requested: We could not find the astronomy handbook or astronomy laptop (neither in the Hab, Science Dome or RAM)

Sol Summary – February 9th

Crew 204 Sol Summary Report 09- FEB – 2019

Sol: 12

Summary Title: Expanding the family

Author’s name: Kunal Naik (Space_Bot)

Mission Status: Nominal

Sol Activity Summary: The time has come for crew 204 to return home. There are mixed feelings in the mind of the crew members. It was a roller coaster ride but ended quite successfully. The crew woke up a bit later than usual, the last warm Martian breakfast with a sip of Coffee. The first thing on the list for the crew was to clean, clean and clean, the Hab as well as all the operation places. Post cleaning was the wait for the upcoming new crew. We were pretty excited to meet the International Emerging Space Leaders Crew (Crew 205). Finally, the wait was over. At noon our Martian family expanded. After the introductions and enjoying lunch together, our crew got to train them on all the basics to survive and efficiently operate around the Hab. The crew 204 handed over the Hab responsibilities to the new crew. Post training, both crews enjoyed burgers together in Hanksville. The crew 205 is an interdisciplinary team, full of enthusiasm and new exciting projects coming up.

Look Ahead Plan: The crew 204 is scheduled to deorbit around 0800 hours and reach Grand Junction for rendezvous with the spacecraft, moving forward to reenter home sweet home.

Anomalies in work: None Encountered

Weather: Sunny

Crew Physical Status: Nominal

EVA’s location: None

Reports to be file: 1) Operations Report 2) Green Hab Report 3) Commander Report 4) Photos

Support Requested: None